# Private Continual Counting of Unbounded Streams

### v1

@@ 3.1 Preliminaries and Notation
The lower-triangular Toeplitz matrix corresponding to a sequence $a_{0},a_{1},\ldots$ is defined as the square matrix whose $i,j$ entry is $a_{i-j}$ when $i\geq j$ and 0 otherwise. For a given sequence, we will denote the corresponding matrix as $\operatorname{LTToep}(a_{0},a_{1},\ldots)$. All matrices we consider will be infinite, but for practical computations we will work with $[A]_{n}$, the leading submatrix formed by the first $n$ rows and columns of $A$. The special matrix $M_{\textit{count}}$ is defined as $\operatorname{LTToep}(1,1,1,\ldots)$.
$\operatorname{LTToep}$ matrices are closed under multiplication, and $\operatorname{LTToep}(S)\operatorname{LTToep}(S^{\prime})=\operatorname{LTToep}(S\ast S^{\prime})$, where $S\ast S^{\prime}$ denotes the convolution of the sequences $S$ and $S^{\prime}$. This means that multiplication of $\operatorname{LTToep}$ matrices is commutative and can be performed in $O(n\log n)$ time using the Fast Fourier Transform (FFT). Given an analytic function $f(z)$ with Taylor coefficients $a_{n}=[z^{n}]f(z)$, we will abuse notation and write $\operatorname{LTToep}(f)$ for $\operatorname{LTToep}(a_{0},a_{1},\ldots)$. With this notation, applying the Cauchy product to their Taylor series gives us that $\operatorname{LTToep}(f)\operatorname{LTToep}(g)=\operatorname{LTToep}(fg)$. We refer the curious reader to the full version of Dvijotham et al. [16] for an excellent and detailed introduction to $\operatorname{LTToep}$ matrices in this context, including proofs of the properties just described.
Throughout, we overload the symbol $\delta$ by using it to refer to both a privacy parameter and the exponent of the $\log\log$ term in Equation 1. This is done to maintain notational consistency with related work, and the meaning should always be clear from context.
@@ 3.2 Main Result
  • Joint Validity: For all $n\in\mathbb{Z}^{+}$, $[L]_{n}[R]_{n}=[M_{\textit{count}}]_{n}$
@@
  • Computability: For any vector $y\in\mathbb{R}^{\infty}$, there exists an unbounded streaming algorithm that at each time step $t$ outputs $(Ly)_{t}$ using $O(t)$ memory and amortized $O(\log t)$ time.
@@
For all $\varepsilon,\delta,\alpha>0$ there exists an unbounded streaming algorithm for continual counting, described in Algorithm 1, which has the same complexity as in Theorem 1 and satisfies $(\varepsilon,\delta)$-DP in the continual release model. At each time step $t$, the algorithm adds Gaussian noise with scale $O\big{(}\log^{1+\alpha}(t)C_{\varepsilon,\delta}\big{)}$, where $C_{\varepsilon,\delta}=O(\varepsilon^{-1}\sqrt{\log(1/\delta)})$ is independent of the input.
We prove Theorem 1 in section 4; Corollary 1 then follows from existing results.  [33, 14].
@@ 3.3 Technical Overview
To understand why these functions in particular are good candidates, consider the special case where $\gamma=\delta=0$. This gives us $R=\operatorname{LTToep}((1-z)^{-1/2})$, which is exactly $M_{\textit{count}}^{1/2}$  [16]. The key issue that makes this sequence inapplicable to unbounded inputs is that it is not square-summable. By Parseval’s theorem, this is equivalent to the statement that the function $f$ is not square-integrable. To fix this, we would like to ‘nudge’ $f$ so that it becomes square-integrable while still being as ‘close’ to the original function as possible. A plausible first attempt in this direction would be to choose $R=\operatorname{LTToep}((1-z)^{-1/2+\alpha})$ for some $\alpha>0$, but this gives $O(t^{\alpha})$ error; the key issue is that $(1-z)^{-\alpha}$ diverges very quickly as $z\to 1^{-}$ even when $\alpha$ is small. To obtain logarithmic overhead, we require functions like $\ln(1/(1-z))$ that diverge more slowly. Multiplying by $f_{2}$ turns out to be sufficient to prove Theorem 1, but we show in section 5 that incorporating $f_{3}$ can further improve our variance by a constant factor. Finally, the inclusion of the $1/z$ and $2/z$ terms in Equation 1 ensures that $f$ is always analytic with $f(0)=1$.
To study the asymptotic growth of $\lVert[L]_{n}\rVert_{2\to\infty}$ and $\lVert[R]_{n}\lVert_{1\to 2}$, we draw on the classic work of Flajolet and Odlyzko [24], whose Theorem 3B provides the following asymptotic expansion for the coefficient of $z^{n}$ in the Taylor series of $f$:
$[z^{n}]f(z;\gamma,\delta)\sim\frac{1}{\sqrt{n\pi}}(\ln n)^{\gamma}(2\ln\ln n)^{\delta}\cdot\Big{(}1+O(\ln^{-1}(n))\Big{)}$
@@
         Sample $y_{s}\sim\mathcal{N}(0,C_{\varepsilon,\delta}^{2}\Delta^{2})$ for $t\leq s\leq 2t-1$
@@
         Compute next $t$ terms of the sequence $Ly$ in $O(t\log t)$ time with an FFT
     Output $S_{t}+(Ly)_{t}$
@@ 4 Proof of Theorem 1
We begin by showing that the column norm of $R$ is finite, or equivalently that $f_{R}\in L^{2}$. Recall that $\lVert R\rVert_{1\to 2}^{2}=\sum_{n=0}^{\infty}([z^{n}]f_{R})^{2}$. By Theorem 3B of Flajolet and Odlyzko [24], we know that $([z^{n}]f_{R})^{2}=O(n^{-1}\log^{-1-2\alpha}(n)\log^{2\delta}(\log n))=O(n^{-1}\log^{-1-\alpha}(n))$, and the series therefore converges by the integral test.
@@
We first consider the case when $\delta\geq 0$. Once, again Theorem 3B of Flajolet and Odlyzko [24] tells us that $([z^{n}]f_{L})^{2}=O(n^{-1}(\ln n)^{1+2\alpha}(\ln\ln n)^{-2\delta})=O(n^{-1}(\ln n)^{1+2\alpha})$. So, for sufficiently large $n>n_{0}$ and some constant $C$, we have $\sum_{m=n_{0}}^{n}([z^{m}]f_{L})^{2}\leq C\sum_{m=n_{0}}^{n}\frac{1}{m}(\ln m)^{1+2\alpha}\leq C\int_{n_{0}-1}^{n}\frac{1}{x}(\ln x)^{1+2\alpha}~{}dx=\frac{C}{2+2\alpha}(\ln(n)^{2+2\alpha}-\ln(n_{0}-1)^{2+2\alpha})$. Therefore, $\sqrt{\sum_{m=0}^{n}([z^{m}]f_{L})^{2}}=O(\ln(n)^{1+\alpha})$ as desired.
The case where $\delta<0$ is similar, except we have $-2\delta>0$ and so it is no longer true that our coefficients are $O(n^{-1}(\ln n)^{1+2\alpha})$. But for any $\alpha^{\prime}>\alpha$, $(\ln\ln n)^{-2\delta}=o((\ln n)^{\alpha^{\prime}-\alpha})$, and so by the same basic argument we can derive an asymptotic error bound of $O(\ln(n)^{1+\alpha+o(1)})$ instead.
@@
We formalize the problem we are trying to solve as a game: first, an adversary secretly fixes some integer $n>0$, as well as an arbitrary vector $y\in\mathbb{R}^{\infty}$ whose elements can be queried in constant time. At each time step $t=1,\ldots,n$, we are asked to output $(\operatorname{LTToep}(f_{L})y)_{t}$, and our goal is to achieve an optimal asymptotic dependence on the unknown value $n$ with respect to the total memory and computation used. We will model computation in units of $M(t)$, the cost of multiplying two polynomials of degree $t$, with the assumption that $M(kt)\sim kM(t)$ for any $k>0$.
@@
Initially, we disregard the challenge of not having access to $n$ and consider the intermediate problem of computing $[\operatorname{LTToep}(f_{L})]_{t}$ for some constant $t$ that we choose for ourselves. As a preview of our eventual doubling trick, we assume that we have access to a pre-computed solution of size $t/2$. From here, we observe that $\operatorname{LTToep}(f(z;-\gamma,-\delta))=\operatorname{LTToep}(f_{1}(z))\operatorname{LTToep}(f_{2}(z;\gamma))\operatorname{LTToep}(f_{3}(z;\delta))$. So, it suffices to compute each of these component matrices in isolation, at which point the full matrix can be computed using $2M(t)$ (or just $M(t)$ if $\delta=0$).
@@
For the first matrix, it is known that $f_{1}(z)=\sum_{m=0}^{\infty}(-z)^{m}{-1/2\choose m}$. This can be used to derive the recurrence relation $[z^{0}]f_{1}=1,~{}[z^{m}]f_{1}=(1-\frac{1}{2m})[z^{m-1}]f_{1}$ presented in prior works  [27, 23]. Using this recurrence, we can compute the coefficients of this matrix in $O(t)$ time and space.
@@
For the second matrix, we have that $[z^{m}]f_{2}(z;1)=\frac{1}{m+1}$, which can be derived from the Taylor expansion of $\ln(1+z)$. This intermediate sequence can also clearly be computed in linear time and space. To account for the power of $-\gamma$, we use the identity $f_{2}(z;-\gamma)=\exp(-\gamma\ln(f_{2}(z;1)))$. To compute the logarithm, we use the fact that $\ln^{\prime}(f_{2}(z;1))=\frac{f_{2}^{\prime}(z;1)}{f_{2}(z^{\prime}1)}$. The coefficients of $f_{2}^{\prime}(z;1)$ are easily computed with term-by-term differentiation, and using the division algorithm of Hanrot and Zimmermann [25] lets us compute the ratio $f_{2}^{\prime}(z;1)/f_{2}(z)$ using $2.25M(t)$. Term-by-term integration then allows us to recover all but the constant term of $\ln(f_{2}(z;1))$ in linear time, which is sufficient because we know $\ln(f_{2}(0;1))=\ln(1)=0$. To compute the exponential, we use the algorithm also presented in  [25] that takes an order $t/2$ approximation of $\exp(h)$ as input and produces an order $t$ approximation using $2.25M(t)$. The same technique can then be applied to the third matrix using an additional $4.5M(t)$ by reusing our earlier computation of $\ln(f_{2}(z;1))$. We can improve this to just $2.25M(t)$ if $\delta=-\gamma$ by computing the quotient $f_{2}(z;1)/f_{3}(z;1)$ before applying the power operation. Doing so also saves us an extra $M(t)$ by removing the need to multiply the second and third matrices at the end.
@@
In total, we require $11M(t)$ to compute $[L]_{t}$, plus $O(t)$ time and space to query the first $t$ coefficients of $y$ and one more $M(t)$ to compute their product. To extend to the case of general $n$, we initialize $t$=2 and double it when we are asked to output the $t+1$st estimate. This gives us a total cost over the entire input of $12\sum_{k=1}^{\lceil\log_{2}n\rceil}M(2^{k})\sim 12\sum_{k=1}^{\lceil\log_{2}n\rceil}2^{-k+1}M(n)\leq 24M(n)$, which can be cut down to $17.5M(n)$ if $\delta=-\gamma$, or $13M(n)$ when $\delta=0$. When polynomial multiplication is implemented with FFTs, we arrive at $O(n\log n)$ time and $O(n)$ space complexity, asymptotically matching the algorithm of Henzinger et al. [27]. But note that because we have to adjust $t$ on the fly, our algorithm requires amortized $O(\log t)$ time at each step, compared to $O(n\log n)$ pre-processing and $O(1)$ time per step for Henzinger et al. [27].
@@ 5 Implementation and Extensions
In this section, we investigate various practical considerations in the implementation of Algorithm 1. We also consider relaxing the assumption that $n$ is completely unknown by giving the algorithm access to (possibly unreliable) side information on its value, and propose an approximate variant (Algorithm 2) which sacrifices some theoretical rigor for a greater than 5×5$\times$5 × improvement in running time for large inputs. Finally, we show that Algorithm 1 can be used as a subcomponent of the hybrid mechanism of Chan et al. [11] to achieve exactly $O(\log^{2}t)$ variance with improved constants compared to their original approach.
@@
There are three natural choices for setting the value of $\delta$, representing different tradeoffs in speed and accuracy. Setting $\delta=0$ is the fastest option computationally, but leads to a sub-optimal rate of growth in error. Choosing $\delta=-\gamma$ is roughly 1.5$\times$ slower, but substantially improves error when $n$ is greater than 1,000 or so. Finally, the special value $\delta=-6\gamma/5$ is notable because it causes the first two Taylor coefficients of both $f_{L}$ and $f_{R}$ to match the function $(1-z)^{-1/2}$ exactly. For a fixed $\gamma$, this gives us the $L$ and $R$ matrices that are closest to $M_{\textit{count}}^{1/2}$, subject to the intuitively reasonable constraint that $L\succeq R$. This option is the most expensive, but empirically leads to improved error for $n$ greater than $2^{20}$, which is shown in Figure 2.
@@
Our basic strategy is to simply pre-compute $Lz$ out to $Cn_{0}$ terms. Provided that the upper bound is true, we recover the same $O(n\log n)$ pre-processing time and $O(1)$ work per-iteration as the $M_{\textit{count}}^{1/2}$ algorithm. Moreover, we suffer no performance cost if the upper bound turns out to be loose because our error at each round depends only on $t$. In contrast, if we used the $M_{\textit{count}}^{1/2}$ algorithm with input $Cn_{0}$, we would be forced to calibrate its noise scale to the sensitivity of the larger matrix $[M_{\textit{count}}^{1/2}]_{Cn_{0}}$. In the unlucky event where it turns out that $n=n_{0}$, this would produce a uniform $\sim 1+\log C/(1+\log n_{0})$ multiplicative scaling of variance across the entire input.
@@
Flajolet and Odlyzko [24] derive the following asymptotic expansion for the Taylor coefficients of $f$, which we used earlier in a weaker form to derive the asymptotic error rate of our mechanism:
$[z^{m}]f(z;\gamma,\delta)\sim\frac{1}{\sqrt{m\pi}}(\ln m)^{\gamma}(\ln\ln m)^{\delta}\Big{(}1+\sum_{k\geq 1}\frac{e_{k}^{(\gamma,\delta)}(\ln\ln m)}{(\ln(m)\ln(\ln(m)))^{k}}\Big{)}$ (2)
where $e_{k}^{(\gamma,\delta)}(x)$ is a polynomial of degree $k$, $e_{k}^{(\gamma,\delta)}(x)=\sqrt{\pi}\frac{d^{k}}{ds^{k}}\big{(}\frac{1}{\Gamma(-s)}\big{)}\big{|}_{s=-1/2}\cdot E_{k}(x)$, and $E_{k}(x)$ is the $k^{th}$ Taylor coefficient of the function $g(u)=(1-xu)^{\gamma}\big{(}1-\frac{1}{x}\ln(1-xu)\big{)}^{\delta}$. While this expression is messy, the cost of computing it is independent of $m$: for fixed approximation order $K$, the derivatives of the reciprocal gamma function can be precomputed and the remaining terms can be calculated in $O(K\log K)$ time using the same techniques described in section 4. This idea naturally suggests Algorithm 2, which switches from exact computations to order $K$ approximations once the relative error of the approximation falls below some threshold $\eta$. We plot relative error as a function of $t,\delta$ and $K$ in Figure 1.
@@
$\gamma,\delta\not\in\{0,1,\ldots\}$, approx. order $K$, tolerance $\eta$, privacy parameter $C_{\varepsilon,\delta}$, sensitivity $\Delta$
@@
         Sample $y_{s}\sim\mathcal{N}(0,(1+\eta)^{2}\Delta^{2}C_{\varepsilon,\delta}^{2})$ for $t\leq s\leq 2t-1$
@@
         $\hat{r}_{t}\leftarrow$ Degree $K$ approximation of $[z^{t-1}]f_{R}$ from Equation 2
@@
         Save computed sequence as $[\hat{R}]_{2t}$
         Compute $[\hat{R}]_{2t}^{-1}[M_{\textit{count}}]_{2t}y$
     Output $(\hat{R}^{-1}M_{\textit{count}}y)_{t}$

### v2

@@ 3.1 Preliminaries and Notation
The lower-triangular Toeplitz matrix corresponding to a sequence $a_{0},a_{1},\ldots$ is defined as the square matrix whose $i,j$ entry is $a_{i-j}$ when $i\geq j$ and 0 otherwise. For a given sequence, we will denote the corresponding matrix as $\operatorname{LTToep}(a_{0},a_{1},\ldots)$. All matrices we consider will be infinite, but for practical computations we will work with $A]_{n}$, the leading submatrix formed by the first $n$ rows and columns of $A$. The special matrix $M_{\textit{count}}$ is defined as $\operatorname{LTToep}(1,1,1,\ldots)$.
$\operatorname{LTToep}$ matrices are closed under multiplication, and $\operatorname{LTToep}(S)\operatorname{LTToep}(S^{\prime})=\operatorname{LTToep}(S\ast S^{\prime})$, where $S\ast S^{\prime}$ denotes the convolution of the sequences $S$ and $S^{\prime}$. This means that multiplication of $\operatorname{LTToep}$ matrices is commutative and can be performed in $O(n\log n)$ time using the Fast Fourier Transform (FFT). Given an analytic function $f(z)$ with Taylor coefficients $a_{n}=[z^{n}]f(z)$, we will abuse notation and write $\operatorname{LTToep}(f)$ for $\operatorname{LTToep}(a_{0},a_{1},\ldots)$. With this notation, applying the Cauchy product to their Taylor series gives us that $\operatorname{LTToep}(f)\operatorname{LTToep}(g)=\operatorname{LTToep}(fg)$. We refer the curious reader to the full version of dvijotham_efficient_2024 for an excellent and detailed introduction to $\operatorname{LTToep}$ matrices in this context, including proofs of the properties just described.
Finally, we note that we reserve the symbol $\delta$ to refer to the exponent of the $\log\log$ term in Equation 1. When we need to refer to the privacy parameter of a DP algorithm, we will use $\delta_{priv}$.
@@ 3.2 Main Result
  • Joint Validity: For all $n\in\mathbb{Z}^{+}$, $L]_{n}[R]_{n}=[M_{\textit{count}}]_{n}$
@@
  • Computability: There exists an unbounded streaming algorithm that at each time step $t$ takes as input $z_{t}$ and outputs $(Lz)_{t}$ using $O(t)$ memory and amortized $O(\log t)$ time
@@
For all $\varepsilon,\delta_{priv},\alpha>0$ there exists an unbounded streaming algorithm for continual counting, described in Algorithm 1, which has the same complexity as in Theorem 1 and satisfies $(\varepsilon,\delta_{priv})$-DP in the continual release model. At each time step $t$, the algorithm adds Gaussian noise with scale $O\big(\log^{1+\alpha}(t)C_{\varepsilon,\delta}\big)$, where $C_{\varepsilon,\delta}=O(\varepsilon^{-1}\sqrt{\log(1/\delta_{priv})})$ is independent of the input.
We prove Theorem 1 in section 4; Corollary 1 then follows from existing results.  [li_matrix_nodate, denisov2022improved].
@@ 3.3 Technical Overview
To understand why these functions in particular are good candidates, consider the special case where $\gamma=\delta=0$. This gives us $R=\operatorname{LTToep}((1-z)^{-1/2})$, which is exactly $M_{\textit{count}}^{1/2}$  [dvijotham_efficient_2024]. The key issue that makes this sequence inapplicable to unbounded inputs is that it is not square-summable. By Parseval’s theorem, this is equivalent to the statement that the function $f$ is not square-integrable. To fix this, we would like to ‘nudge’ $f$ so that it becomes square-integrable while still being as ‘close’ to the original function as possible. A plausible first attempt in this direction would be to choose $R=\operatorname{LTToep}((1-z)^{-1/2+\alpha})$ for some $\alpha>0$, but this gives $O(t^{\alpha})$ error; the key issue is that $(1-z)^{-\alpha}$ diverges very quickly as $z\to 1^{-}$ even when $\alpha$ is small. To obtain logarithmic overhead, we require functions like $\ln(1/(1-z))$ that diverge more slowly. Multiplying by $f_{2}$ turns out to be sufficient to prove Theorem 1, but we show in section 5 that incorporating $f_{3}$ can further improve our variance by a constant factor. Finally, the inclusion of the $1/z$ and $2/z$ terms in Equation 1 eliminates the 0 of $\ln(1/(1-z))$ at $z=0$, which ensures that $f$ is always analytic with $f(0)=1$.
To study the asymptotic growth of $\lVert[L]_{n}\rVert_{2\to\infty}$ and $\lVert[R]_{n}\lVert_{1\to 2}$, we draw on the classic work of flajolet1990singularity, whose Theorem 3B (reproduced fully in Theorem 2) provides the following asymptotic expansion for the coefficient of $z^{n}$ in the Taylor series of $f$:
$z^{n}]f(z;\gamma,\delta)\sim\frac{1}{\sqrt{n\pi}}(\ln n)^{\gamma}(2\ln\ln n)^{\delta}\cdot\Big(1+O(\ln^{-1}(n))\Big)$
@@
   Sample $z_{s}\sim\mathcal{N}(0,C_{\varepsilon,\delta}^{2}\Delta^{2})$ for $t\leq s\leq 2t-1$
@@
   Compute next $t$ terms of the sequence LzLz in $O(t\log t)$ time with an FFT
  Output $S_{t}+(Lz)_{t}$
@@ 4 Proof of Theorem 1
We begin by showing that the column norm of $R$ is finite, or equivalently that $f_{R}\in L^{2}$. Recall that $\lVert R\rVert_{1\to 2}^{2}=\sum_{n=0}^{\infty}([z^{n}]f_{R})^{2}$. By Theorem 3B of flajolet1990singularity, we know that $([z^{n}]f_{R})^{2}=O(n^{-1}\log^{-1-2\alpha}(n)\log^{2\delta}(\log n))=O(n^{-1}\log^{-1-\alpha}(n))$, and the series therefore converges by the integral test.
@@
We first consider the case when $\delta\geq 0$. Once, again Theorem 3B of flajolet1990singularity tells us that $([z^{n}]f_{L})^{2}=O(n^{-1}(\ln n)^{1+2\alpha}(\ln\ln n)^{-2\delta})=O(n^{-1}(\ln n)^{1+2\alpha})$. So, by the definition of big-$O$, there are constants $n_{0},C$ such that for all $n>n_{0}$, we have $\sum_{m=n_{0}}^{n}([z^{m}]f_{L})^{2}\leq C\sum_{m=n_{0}}^{n}\frac{1}{m}(\ln m)^{1+2\alpha}\leq C\int_{n_{0}-1}^{n}\frac{1}{x}(\ln x)^{1+2\alpha}~dx=\frac{C}{2+2\alpha}(\ln(n)^{2+2\alpha}-\ln(n_{0}-1)^{2+2\alpha})$. Therefore, $\sqrt{\sum_{m=0}^{n}([z^{m}]f_{L})^{2}}=O(\ln(n)^{1+\alpha})$ as desired.
The case where $\delta<0$ is similar, except we have $2\delta>0$ and so it is no longer true that our coefficients are $O(n^{-1}(\ln n)^{1+2\alpha})$. But for any $\alpha^{\prime}>\alpha$, $(\ln\ln n)^{-2\delta}=o((\ln n)^{\alpha^{\prime}-\alpha})$, and so by the same basic argument we can derive an asymptotic error bound of $O(\ln(n)^{1+\alpha+o(1)})$ instead.
@@
We formalize the problem we are trying to solve as a game: first, an adversary secretly fixes some integer $n>0$. Then, at each time step $t=1,\ldots,n$, our algorithm is required to output $(\operatorname{LTToep}(f_{L})z)_{t}$, where $z\sim\mathcal{N}(0,I_{n})$. The algorithm’s goal is to achieve an optimal asymptotic dependence on the unknown value $n$ with respect to the total memory and computation used. We will model computation in units of $M(t)$, the cost of multiplying two polynomials of degree $t$. Because polynomial multiplication can be implemented through divide-and-conquer algorithms, we assume that for any $k>0$, $M(kt)\sim kM(t)$ asymptotically as $t\to\infty$.
@@
Initially, we disregard the challenge of not having access to $n$ and consider the intermediate problem of computing $\operatorname{LTToep}(f_{L})]_{t}$ for some constant $t$ that we choose ourselves. As a preview of our eventual strategy for handling unbounded inputs, we assume that we have access to a pre-computed solution of size $t/2$. From here, we observe that $\operatorname{LTToep}{f(z;-\gamma,-\delta)}=\operatorname{LTToep}(f_{1}(z))\operatorname{LTToep}(f_{2}(z;\gamma))\operatorname{LTToep}(f_{3}(z;\delta))$. So, it suffices to compute each of these matrices in isolation. The final product can then be computed using $2M(t)$ (or just $M(t)$ if $\delta=0$).
@@
To compute $\operatorname{LTToep}(f_{1}(z))$, we use the fact that $f_{1}=\sum_{m=0}^{\infty}(-x)^{m}{-1/2\choose m}$. This can be used to derive the recurrence relation $z^{0}]f_{1}=1,~[z^{m}]f_{1}=(1-\frac{1}{2m})[z^{m-1}]f_{1}$ presented in prior works  [henzinger_almost_2023, fichtenberger2023constant]. This recurrence lets us compute the coefficients of this matrix in $O(t)$ time and space.
@@
To compute the cofficients of $\operatorname{LTToep}(f_{2}(z;\gamma)$, we begin with the fact that $z^{m}]f_{2}(z;1)=\frac{1}{m+1}$, which can be derived from the Taylor expansion of $\ln(1+z)$. This intermediate sequence can also clearly be computed in linear time and space. To account for the power of $\gamma$, we use the identity $f_{2}(z;-\gamma)=\exp(-\gamma\ln(f_{2}(z;1)))$. To compute $\ln(f_{2}(z;1))$, we use the fact that $\ln^{\prime}(f_{2}(z;1))=\frac{f_{2}^{\prime}(z;1)}{f_{2}(z^{\prime}1)}$. The coefficients of $f_{2}^{\prime}(z;1)$ can be directly computed from the coefficients of $f_{2}(z;1)$ through term-by-term differentiation, and using the division algorithm of hanrot2004newton lets us compute the ratio $f_{2}^{\prime}(z;1)/f_{2}(z)$ using $2.25M(t)$. Term-by-term integration then allows us to recover all but the constant term of $\ln(f_{2}(z;1))$ in linear time, but this is sufficient because we know $\ln(f_{2}(0;1))=\ln(1)=0$.
Having shown that we can efficiently compute the Taylor coefficients of $\gamma\ln(f_{2}(z;1))$, it remains to find the coefficients of $f_{2}(z;-\gamma)=\exp(-\gamma\ln(f_{2}(z;1))$. We achieve this by using the algorithm also presented in  [hanrot2004newton] that takes an order $t/2$ approximation of $\exp(h)$ as input and produces an order $t$ approximation using $2.25M(t)$.
Finally, we can rewrite $f_{3}(z;\delta)$ as $(\frac{2}{z}\ln(f_{2}(z;1))^{\delta}$. The allows us to reuse our earlier computation of the Taylor coefficients of $\ln(f_{2}(z;1))$. Dividing by $z$ corresponds to a simple shift of the coefficients. All that remains is to account for the power of $\delta$, which can be done in $4.5M(t)$ using the same technique described above for $\gamma$. We can improve this to just $2.25M(t)$ if $\delta=-\gamma$ by computing the quotient $f_{2}(z;1)/f_{3}(z;1)$ before applying the power operation. Doing so also saves us an extra $M(t)$ by removing the need to multiply $\operatorname{LTToep}(f_{2}(z;\gamma)$ and $\operatorname{LTToep}(f_{3}(z;\delta)$ at the end.
@@
In total, we require $11M(t)$ to compute $L]_{t}$, plus one more $M(t)$ to compute $L]_{t}z$. To extend to the case of general $n$, we initialize tt=2 and double it when we are asked to output the t+1t+1st estimate. This gives us a total cost over the entire input of $12\sum_{k=1}^{\lceil\log_{2}n\rceil}M(2^{k})\sim 12\sum_{k=1}^{\lceil\log_{2}n\rceil}2^{-k+1}M(n)\leq 24M(n)$, which can be cut down to $17.5M(n)$ if $\delta=-\gamma$, or $13M(n)$ when $\delta=0$. When polynomial multiplication is implemented with FFTs, we arrive at $O(n\log n)$ time and $O(n)$ space complexity, asymptotically matching the algorithm of henzinger_almost_2023. But note that because we have to adjust $t$ on the fly, our algorithm requires amortized $O(\log t)$ time at each step, compared to $O(n\log n)$ pre-processing and $O(1)$ time per step for henzinger_almost_2023.
@@ 5 Implementation and Extensions
In this section, we investigate various practical considerations in the implementation of Algorithm 1. We also consider relaxing the assumption that $n$ is completely unknown by giving the algorithm access to (possibly unreliable) side information on its value, and propose an approximate variant (Algorithm 2) which sacrifices some theoretical rigor for a greater than $5\times$ improvement in running time for large inputs. Finally, we show that Algorithm 1 can be used as a subcomponent of the hybrid mechanism of chan2011private to achieve exactly $O(\log^{2}t)$ variance with improved constants compared to their original approach.  [^1] Our code is available at https://github.com/ben-jacobsen/central-dpolo/
@@
[…]
@@
Our basic strategy is to simply pre-compute $Lz$ out to $Cn_{0}$ terms. Provided that the upper bound is true, we recover the same $O(n\log n)$ pre-processing time and $O(1)$ work per-iteration as the $M_{\textit{count}}^{1/2}$ algorithm. Moreover, we suffer no performance cost if the upper bound turns out to be loose because our error at each round depends only on $t$. In contrast, if we used the $M_{\textit{count}}^{1/2}$ algorithm with input $Cn_{0}$, we would be forced to calibrate its noise scale to the sensitivity of the larger matrix $M_{\textit{count}}^{1/2}]_{Cn_{0}}$. In the unlucky event where it turns out that $n=n_{0}$, this would produce a uniform $\sim 1+\log C/(1+\log n_{0})$ multiplicative scaling of variance across the entire input.
@@
flajolet1990singularity derive the following asymptotic expansion for the Taylor coefficients of $f$, which we used earlier in a weaker form to derive the asymptotic error rate of our mechanism:
Theorem 2 (Paraphrased from flajolet1990singularity Theorem 3B).
Let $\gamma,$ and $\delta$ be complex numbers not in $\{0,1,2,\ldots\}$. Then the Taylor coefficients of $f(z;\gamma,\delta)$ satisfy:
$z^{m}]f(z;\gamma,\delta)\sim\frac{1}{\sqrt{m\pi}}(\ln m)^{\gamma}(\ln\ln m)^{\delta}\Big(1+\sum_{k\geq 1}\frac{e_{k}^{(\gamma,\delta)}(\ln\ln m)}{(\ln(m)\ln(\ln(m)))^{k}}\Big)$ (2)
where $e_{k}^{(\gamma,\delta)}(x)$ is a polynomial of degree $k$:
$e_{k}^{(\gamma,\delta)}(x)=\sqrt{\pi}\frac{d^{k}}{ds^{k}}\big(\frac{1}{\Gamma(-s)}\big)\big|_{s=-1/2}\cdot E_{k}(x)$
and $E_{k}(x)$ is the $k^{th}$ Taylor coefficient of the function $g(u)=(1-xu)^{\gamma}\big(1-\frac{1}{x}\ln(1-xu)\big)^{\delta}$.
While this expression is messy, the cost of computing it is independent of $m$: for fixed approximation order $K$, the derivatives of the reciprocal gamma function can be precomputed and the remaining terms can be calculated in $O(K\log K)$ time using the same techniques described in section 4. This idea naturally suggests Algorithm 2, which switches from exact computations to order $K$ approximations once the relative error of the approximation falls below some threshold $\eta$. We plot relative error as a function of $t,\delta$ and $K$ in Figure 1.
@@
Matrix parameters $\gamma,\delta\not\in\{0,1,\ldots\}$, order $K$, error tolerance $\eta$, privacy parameter $C_{\varepsilon,\delta}$
@@
   Sample $z_{s}\sim\mathcal{N}(0,(1+\eta)^{2}C_{\varepsilon,\delta}^{2})$ for $t\leq s\leq 2t-1$
@@
   $\hat{r}_{t}\leftarrow$ Degree $K$ approximation of $z^{t-1}]f_{R}$ from Equation 2
@@
   Save computed sequence as $\hat{R}]_{2t}$
   Compute $\hat{R}]_{2t}^{-1}[M_{\textit{count}}]_{2t}z$
  Output $(M_{\textit{count}}\hat{R}^{-1}z)_{t}$